The measure of the area of a rectangle is 6x2+13x-5. If the demensions of nthe rectangle are represented by polynomials, find the demensions of the rectangle and then find its perimeter.
Answer by ptaylor(2198)

Area of rectangle=L*W=6x^2+13x-5
6x^2+13x-5=(3x-1)(2x+5)
so, L*W=(3x-1)*(2x+5)--------------dimensions
Perimeter=2L+2W=2(3x-1)+2(2x+5)=6x-2+4x+10=10x+8

Does this help???------ptaylor